Under normal conditions the human heart converts about 13.3 J of chemical energy per second into 1.33 W of mechanical power as it pumps blood throughout the body.


(a)

Determine the number of Calories required to power the heart for one day, given that 1 Calorie equals 4186 J.

__ Cal


(b)

Metabolizing 1 kg of fat can release about 9000 Calories of energy. What mass of metabolized fat (in kg) would power the heart for one day?

__ kg


1
Expert's answer
2021-09-23T15:48:38-0400

In one day "= 13.3 \times 86400 \;s = 1149120\; J"

"= \frac{1149128}{4186} \\

= 274.52 \;cal"

(b) 1 kg = 9000 cal

"274.52 \;cal = \frac{274.52}{9000} \;kg \\

= 0.0305 \;kg"